OpenStudy (anonymous):

What is the equation of the line which includes points (2, 6) and (0, 0)? A. y = -3x + 1 B. y = -3x C. y = 3x D. y = 3x + 1

OpenStudy (anonymous):

plug the points in an check...

OpenStudy (anonymous):

find the slope 1st \[m=\frac{ y2-y1 }{ x2-x1 }\] y = mx + b

OpenStudy (anonymous):

im here, did you miss me that much lol y = mx + b you found m, now choose one point and plug it for x and y

OpenStudy (anonymous):

@charlotteakina

OpenStudy (anonymous):

either or give you same answer but i will choose (0,0)

OpenStudy (anonymous):

y = mx + b 0 = 3*0 + b b = 0

OpenStudy (anonymous):

now you try the other point

OpenStudy (anonymous):

should get same answer

OpenStudy (anonymous):

Its 6=3(2)+b

OpenStudy (anonymous):

6=6+b find b

OpenStudy (anonymous):

OK since you found that the slope(m) is positive, you have either C or D

OpenStudy (anonymous):

But the b which is the intercept is 0 so its y=3x+0 or y=3x which is C

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

C. y = 3x + 0